Need computation for this :)

Mr. Jojo, married, supporting his 6 children (2 of which are gainfully employed) had the following data for taxable year 2019:
Phil.                Abroad
Business income                     P1,000,000           $20,000
Professional income                    400,000            10,000
Salaries                                          200,000           
Business and professional
Expenses 250,000               8,000
Income tax paid                             200,000                 -

Note: $1=P50

1. If he is a non-resident alien not engage in trade or business, disregard professional

Income and salaries, his income tax due? _________
Answer #1

Assuming all data of Phil. Is in P Currency and all data of Abroad is in $

First convert Data of Phil into dollars

Business Income = 1000000 / 50 = $20000

Professional Income = 400000 / 50 =$ 8000

Salaries = 200000 / 50 = $4000

Income Tax Paid = 200000 / 50 = $4000

Business and Profession Expenses = 250000 / 50 = $5000

Now combined Gross Income from Business = 20000 + 20000 =40000

Next combined Gross Professional Income = 8000 + 10000 = 18000

Net Income from Business & Profession = Gross Income from Business & Profession Less expnses

                                                                                = 40000 + 18000 – 5000 – 8000

                                                                                = 45000

Income from Salaries                                     = 4000

Net Income                                                        = 45000 + 4000

                                                                                = 49000

Tax on Net Income                                          = 49000 * 30%

                                                                                = 14700

Less Tax already paid                                      = 4000

Net Tax due                                                       = $10700